### v3

@@ Abstract
Experimental techniques for the characterisation of three-dimensional (3D) magnetic spin structures are required to advance the performance of nanoscale magnetic technologies. However, as component dimensions approach the nanometre range, it becomes ever more challenging to analyse 3D magnetic configurations quantitatively with the required spatial resolution and sensitivity. Here, we use off-axis electron holography and model-based iterative reconstruction to reconstruct the 3D magnetisation distribution in an exemplary nanostructure comprising an L-shaped ferromagnetic cobalt nanowire fabricated using focused electron beam induced deposition. Our approach involves using off-axis electron holography to record tomographic tilt series of electron holograms, which are analysed to reconstruct electron optical magnetic phase shifts about two axes with tilts of up to ±$\pm 60\degree$. A 3D magnetisation vector field that provides a best fit to the tomographic phase measurements is then reconstructed, revealing multiple magnetic domains in the nanowire. The reconstructed magnetisation is shown to be accurate for magnetic domains that are larger than approximately 50 nm. Higher spatial resolution and improved signal to noise can be achieved in the future by using more specialised electron microscopes, improved reconstruction algorithms, and automation of data acquisition and analysis.
@@ 1 Introduction
where $\vec{M}$ is the magnetisation vector field describing magnetic moment per unit volume, $\vec{B}$ is the magnetic induction and $\mu_{0}$ is the vacuum permeability. If no external magnetic fields, conduction currents or displacement currents are present, then $\vec{H}$ is the demagnetising field, which is defined by the distribution of $\vec{M}$. X-ray imaging currently allows 3D reconstruction of $\vec{M}$ with sub-50-nm resolution using magnetic laminography (Donnelly et al., 2020) and sub-70-nm resolution using Fourier transform holography (Donnelly et al., 2017; Di Pietro Martínez et al., 2023). Theoretical predictions estimate a resolution for X-ray laminography down to 20 nm (Donnelly et al., 2022). In comparison, phase contrast techniques in the transmission electron microscope (TEM) cannot be used to reconstruct $\vec{M}$ directly, but have been used to measure $\vec{B}$ with a 3D resolution of 10 nm using electron holographic vector field tomography (EH-VFT) (Phatak et al., 2008, 2010; Wolf et al., 2019). Micromagnetic simulations are often used to interpret TEM measurements of $\vec{B}$ (Donahue and McMichael, 1997; Lyu et al., 2024). Here, we use the alternative approach of applying model-based iterative reconstruction (MBIR) to reconstruct a magnetisation distribution $\vec{M}$ from experimental measurements that are sensitive to $\vec{B}$ rather than $\vec{M}$. We discuss how magnetic and electrostatic signals obtained from electron phase measurements can be used to improve the accuracy of the reconstruction, as well as practical limitations to resolution and accuracy.
@@ 1.2 Model-based iterative reconstruction of $\vec{M}$
The theoretical development of MBIR for the tomographic reconstruction of $\vec{M}$ from EH-VFT measurements has been described in detail elsewhere (Caron, 2018a). The most common phase contrast techniques for magnetic imaging in the TEM (Lubk and Zweck, 2015; Kohn et al., 2016; Krajnak et al., 2016; You et al., 2023; Cui et al., 2024), which are often collectively termed Lorentz microscopy, all rely fundamentally on an interpretation of the Aharonov-Bohm effect (Aharonov and Bohm, 1959), which describes how the phase of an electron wave is affected by electromagnetic potentials. The interaction of an electron wave travelling with a magnetic induction field $\vec{B}$ results in an electron phase shift $\varphi_{mag}$, where
@@
The workflow for the reconstruction of $\vec{M}$ (Fig. 1) includes: 1) Processing and alignment of experimental data to construct a 3D geometrical model of the sample; 2) MBIR reconstruction of the $\vec{M}$ distribution that best matches the experimental data; 3) Optimal estimation diagnostics for evaluating reconstruction errors. The theory of applying MBIR for $\vec{M}$ reconstruction has been developed and tested extensively using simulated datasets (Caron, 2018b). The present work is an experimental development, which addresses several sources of error, such as those caused by sample alignment (Diehle et al., 2021), sample drift, and surface damage. It also builds on other previous work on the 3D reconstruction of $\vec{M}$ from experimental electron microscopy data (Mohan et al., 2018).
@@ 1.3 Null spaces in electron phase measurement
In Lorentz microscopy, some configurations of $\vec{M}$ (Mansuripur, 1991) contribute no signal to experimental measurements made in projection and can be described as null spaces, where $\vec{M}$ may be reconstructed falsely as being close to zero (Caron, 2018c). In such projections, the vector components of $\vec{M}$ can sum to zero or have non-zero divergence of $\vec{M}$. In some cases, tiltng the sample can reveal the hidden components. However, divergent components can remain undetectable even when tilting of the sample is used. As a result, some configurations of $\vec{M}$, such as Néel-type magnetic domain walls, cannot be reconstructed fully without making using of additional constraints during reconstruction. Examples of the application of MBIR to such states are discussed in the Supplementary Material [S2]. It should be noted that null spaces are not present in the experimental reconstruction described in this paper.
@@ 2.2 Electron holographic vector field tomography
Tilt series of off-axis electron holograms were recorded at 10°$\degree$ tilt increments over two arcs. The effective maximum tilt angles were limited to between -60°$\degree$ and 30°$\degree$ and to between -60°$\degree$ and 0°$\degree$, as the sample was shadowed by the holder and the substrate. The second tilt series was recorded to reduce missing wedge artefacts and to prevent projection-based null spaces (Caron, 2018c). Simulations of samples comprising uniformly-magnetised nanowires and vortex configurations have shown that the use of smaller tilt increments has a negligible impact on reconstruction quality (Caron, 2018b). Furthermore, simulations presented in the Supplementary Material [S1] show that if a 10% difference between the volumes of the sample and its geometrical model is assumed then the difference between the reconstruction of $\vec{M}$ and the ground truth is on average less than 10% per voxel. In order to separate the electrostatic and magnetic contributions to the phase images, off-axis electron holograms were recorded before and after turning the sample by 180°$\degree$. This approach changes the sign of the magnetic contribution to the total phase shift, based on the expression (Aharonov and Bohm, 1959; Weyland and Midgley, 2016),
@@
where $\varphi_{el}(V)$ and $\varphi_{mag}(\vec{A})$ are the electrostatic and magnetic contributions to the phase shift, respectively, $\nu$ is the relativistic electron speed and $\text{d}z$ and $\text{d}\vec{z}$ are scalar and vector elements of the electron path, respectively. For a pair of such reversed holograms, half of the sum and half of the difference of the measured phase images represent $\varphi_{el}$ and $\varphi_{mag}$, respectively. Unfortunately, physically turning the sample over by 180°$\degree$ introduces misalignments between the phase images. These misalignments were corrected by using an automated Python script to identify matching features at the sample edges (Walt et al., 2014; Peña et al., 2019; Paterson et al., 2020). Linear distortions caused by the misalignment of the electron optical system of the TEM were corrected by using an affine transformation to provide an optimal match between the edges of each phase image pair, as shown in Fig. 4c. In the present study, the images were found to be elongated by 2% in a direction that forms a 75°$\degree$ angle with the tilt axis by tilting a calibration sample by 180°$\degree$ and tracking fiducial markers. By applying the affine transformation, phase separation artefacts were reduced to below the noise level, such that they were not detectable in line profiles. This approach ensures a consistent separation of $\varphi_{el}$ and $\varphi_{mag}$, as shown in the inset to Fig. 4b. All necessary software packages are listed in Supplementary Material [S3], and examples of intermediate steps of the hologram-to-phase calculation are shown in Supplementary Material S4.
@@ 2.3 Alignment of phase images
As each $\varphi_{mag}$ image can include pixels at which the hologram-to-phase calculation resulted in reconstruction or misalignment errors, a confidence mask was used to identify the areas in each phase image that were suitable for reconstruction. The $\varphi_{el}$ image was first flattened to remove background ramps resulting from effects such as electron-beam-induced charging and the perturbed reference wave. A threshold was then used to define a mask that identified where the sample is located at each tilt angle. If elemental mapping was performed, as shown in Fig. 3e, then the threshold could be fine-tuned to include only ferromagnetic material in the mask. These binary masks were used for aligning the phase images in each tilt series. Since the present phase images have a spatial resolution of no better than 6 nm due to the chosen hologram fringe spacing, tomographic reconstruction was found to be highly sensitive to small image misalignments. Furthermore, the sample orientation was not known precisely during acquisition, as only relative tilt angles between phase images $\Delta\theta_{i}$ were measured. The projection positions shown in Fig. 5a were therefore refined by using an axial-symmetry-based implementation of the common lines method (Penczek et al., 1996). As illustrated in Fig. 5b, at every sample tilt angle $\Delta\theta_{i}$ there is a unique projection orientation $\alpha_{i}$, and the two are related via geometrical parameters: $\phi$ is the observed detector rotation relative to the sample tilt axis; $\theta_{0}$ is the starting tilt of the sample; $\alpha_{0}$ is the sample orientation at 0°$\degree$ tilt. $\alpha_{i}$ can be measured by finding the sample symmetry axis from the electrostatic phase maps, as shown in Fig. 5c. The equation
@@
Figure 5: Alignment of tilt series of phase images and generation of a 3D geometrical model of the sample. (a) Schematic diagram showing the angular position and sample outline of each phase image used in the reconstruction. The direction of the electron beam is shown in red. The arcs of rotation of the sample are shown in black. (b) Rotations that define the projections for a conically-symmetrical element, showing the sample tilt $\Delta\theta_{i}$, the projection direction $\alpha_{i}$ and the rotation between the detector and the tilt axis $\phi$. (c) The centre of an axially-symmetrical part of the sample can be identified in each projection and used to align the projections. (d) 3D tomographic reconstruction of the masks in the aligned projections performed to generate a model representing the sample geometry. A 10.2 nm voxel size was used to increase the computational speed.
@@ 3.1 Uniform region reconstruction
As described in the preceding Section, EH-VFT and MBIR were used to record and reconstruct the 3D distribution of $\vec{M}_{rec}$ in a sample comprising two intersecting nanowires. Fig. 3e shows that the nanowires contain a mixture of cobalt and organic compounds, which transitions from a 60% cobalt content in the core to a pure carbon shell at the surface. Since the cobalt is dispersed within the carbon matrix in nanocrystalline form (Pablo-Navarro et al., 2018) and bulk cobalt has $\mu_{0}M_{s}~{}=~{}1.75$ T (Pablo-Navarro et al., 2018), the nanowires should exhibit $\mu_{0}M_{s}~{}<~{}1.1$ T, as observed for the nanowire distribution in Fig. 10. The few outliers with $\mu_{0}M_{s}~{}>~{}1.1$ T are likely to be artefacts at the interface between the magnetic material and the vacuum region, where magnetic voxels that are designated as non-magnetic force the reconstruction to increase the value of $M_{s}$ locally. In other parts of the sample, surface voxels contain non-magnetic material where $M_{s}$ is close to zero. Surface reconstruction artefacts are present in part because missing wedge correction for the geometrical model is not perfect, as discussed in the Supplementary Material [S1]. The threshold for creating the 3D geometrical model shown in Fig. 5d was refined after the first reconstruction, so that similar amounts of over- and underestimating surface artefacts were present. Such surface artefacts are also identifiable for the full sample, as voxels with $\mu_{0}M_{s}~{}>~{}1.75$ T, as shown in the histogram in Fig. 10. Such errors could be reduced in the future by using smaller voxels and improved tomographic alignment and distortion removal algorithms. Both nanowires have single magnetic domain configurations at a distance of 200 nm from the intersection, with $\mu_{0}M_{s}~{}=~{}0.7\pm 0.2$ T in the left nanowire and $\mu_{0}M_{s}~{}=~{}0.5\pm 0.2$ T in the right nanowire. $M_{s}$ has previously been measured for varying cobalt contents of FEBID nanowires (Teresa et al., 2016; Pablo-Navarro et al., 2018), but the measurements were corrected to exclude the non-magnetic shell. The latter correction can be replicated by assuming that in Fig. 3e the compositional line profile corresponds to a circular-cross-section nanowire with a 100-nm-diameter magnetic core, resulting in an estimated average cobalt content of 61% with 4% standard deviation, which would correspond to $\mu_{0}M_{s}~{}=~{}0.5\pm 0.1$ T. The reconstructed value of $\vec{M}_{rec}$ is consistent with the predicted values, but shows a higher standard deviation due to surface artefacts.

### v4

@@ Abstract
Experimental techniques for the characterisation of three-dimensional (3D) magnetic spin structures are required to advance the performance of nanoscale magnetic technologies. However, as component dimensions approach the nanometre range, it becomes ever more challenging to analyse 3D magnetic configurations quantitatively with the required spatial resolution and sensitivity. Here, we use off-axis electron holography and model-based iterative reconstruction to reconstruct the 3D magnetisation distribution in an exemplary nanostructure comprising an L-shaped ferromagnetic cobalt nanowire fabricated using focused electron beam induced deposition. Our approach involves using off-axis electron holography to record tomographic tilt series of electron holograms, which are analysed to reconstruct electron optical magnetic phase shifts about two axes with tilts of up to ±$\pm 60\degree$. A 3D magnetisation vector field that provides the best fit to the tomographic phase measurements is then reconstructed, revealing multiple magnetic domains in the nanowire. The reconstructed magnetisation is shown to be accurate for magnetic domains that are larger than approximately 50 nm. Higher spatial resolution and improved signal-to-noise can be achieved in the future by using more specialised electron microscopes, improved reconstruction algorithms, and automation of data acquisition and analysis.
@@ 1 Introduction
where $\vec{M}$ is the magnetisation vector field describing magnetic moment per unit volume, $\vec{B}$ is the magnetic induction and $\mu_{0}$ is the vacuum permeability. If no external magnetic fields, conduction currents or displacement currents are present, then $\vec{H}$ is the demagnetising field, which is defined by the distribution of $\vec{M}$. It should be noted that, in general, $\vec{B}$ permeates all of space, while $\vec{M}$ is local to the sample and has a more compact description than $\vec{B}$. X-ray imaging currently allows 3D reconstruction of $\vec{M}$ with sub-50-nm resolution using magnetic laminography (Donnelly et al., 2020) and sub-70-nm resolution using Fourier transform holography (Donnelly et al., 2017; Di Pietro Martínez et al., 2023). Theoretical predictions estimate a resolution for X-ray laminography down to 20 nm (Donnelly et al., 2022). In comparison, phase contrast techniques in the transmission electron microscope (TEM) cannot be used to reconstruct $\vec{M}$ directly, but have been used to measure $\vec{B}$ with a 3D resolution of 10 nm using electron holographic vector field tomography (EH-VFT) (Phatak et al., 2008, 2010; Wolf et al., 2019). Micromagnetic simulations have also been used to interpret TEM measurements of $\vec{B}$ (Donahue and McMichael, 1997; Lyu et al., 2024). Here, we use the alternative approach of applying model-based iterative reconstruction (MBIR) to reconstruct a magnetisation distribution $\vec{M}$ from experimental TEM measurements that are sensitive to $\vec{B}$ rather than $\vec{M}$. We discuss how magnetic and electrostatic signals obtained from electron phase measurements can be used to improve the reconstruction, as well as practical limitations to resolution and accuracy. Whereas most previous TEM studies have used EH-VFT to reconstruct $\vec{B}$ in 3D, we demonstrate the reconstruction of $\vec{M}$ in 3D by applying MBIR to an EH-VFT dataset.
@@ 1.2 Model-based iterative reconstruction of $\vec{M}$
The theoretical development of MBIR for the tomographic reconstruction of $\vec{M}$ from EH-VFT measurements has been described in detail elsewhere (Caron, 2018a). The most common phase contrast techniques for magnetic imaging in the TEM (Lubk and Zweck, 2015; Kohn et al., 2016; Krajnak et al., 2016; You et al., 2023; Cui et al., 2024), which are often collectively termed Lorentz microscopy, all rely fundamentally on an interpretation of the Aharonov-Bohm effect (Aharonov and Bohm, 1959), which describes how the phase of an electron wave is affected by electromagnetic potentials. The interaction of an electron wave with a magnetic induction field $\vec{B}$ results in an electron phase shift $\varphi_{mag}$, where
@@
The workflow for the reconstruction of $\vec{M}$ (Fig. 1) includes: 1) Processing and alignment of experimental data to construct a 3D geometrical model of the sample; 2) MBIR reconstruction of the $\vec{M}$ distribution that best matches the experimental data; 3) Optimal estimation diagnostics for evaluating reconstruction errors. The theory of applying MBIR for $\vec{M}$ reconstruction has been developed and tested extensively using simulated datasets (Caron, 2018b). The present work is an experimental development, which addresses several sources of error, such as those caused by sample alignment (Diehle et al., 2021), sample drift, and surface damage. It also builds on other previous work on the 3D reconstruction of $\vec{M}$ from experimental electron microscopy data (Aditya Mohan et al., 2018).
@@ 1.3 Null spaces in electron phase measurement
In Lorentz microscopy, some configurations of $\vec{M}$ (Mansuripur, 1991) contribute no signal to experimental measurements made in projection and can be described as null spaces, where $\vec{M}$ may be reconstructed falsely as being close to zero (Caron, 2018c). In such projections, the vector components of $\vec{M}$ can sum to zero or have non-zero divergence of $\vec{M}$. In some cases, tiltng the sample can reveal the hidden components. However, divergent components can remain undetectable even when tilting of the sample is used. As a result, some configurations of $\vec{M}$, such as Néel-type magnetic domain walls, cannot be reconstructed fully because they contain some states of $\vec{M}$ that do not create a detectable TEM signal in any projection. It has been observed experimentally (McVitie et al., 2018) and proven mathematically (Caron, 2018c) that Néel-type magnetic domain walls in both out-of-plane and in-plane magnetised films include eigenstates of the $\vec{M}$ distribution that do not result in electron phase shift. Such states cannot be reconstructed without making use of additional constraints, and examples of the application of MBIR to such states are discussed in the Supplementary Material [S2]. It should be noted that null spaces are not present in the experimental reconstruction described in this paper.
@@ 2.2 Electron holographic vector field tomography
Tilt series of off-axis electron holograms were recorded at 10°$\degree$ tilt increments over two orthogonal arcs. The effective maximum tilt angles were limited to between -60°$\degree$ and 30°$\degree$ and to between -60°$\degree$ and 0°$\degree$, as the sample was shadowed by the holder and the substrate. For each arc, holograms were recorded at the same tilt angles before and after turning the sample over, in order to enable phase shift separation, which is described below. The second arc was recorded to reduce missing wedge artefacts and to prevent projection-based null spaces (Caron, 2018c). Simulations of samples comprising uniformly-magnetised nanowires and vortex configurations have shown that the use of smaller tilt increments has a negligible impact on reconstruction quality (Caron, 2018b). Furthermore, simulations presented in the Supplementary Material [S1] show that if a 10% difference between the volumes of the sample and its geometrical model is assumed then the difference between the reconstruction of $\vec{M}$ and the ground truth is on average less than 10% per voxel. In order to separate the electrostatic and magnetic contributions to the phase images, off-axis electron holograms were recorded before and after turning the sample by 180°$\degree$. This approach changes the sign of the magnetic contribution to the total phase shift, based on the expression (Aharonov and Bohm, 1959; Weyland and Midgley, 2016),
@@
where $\varphi_{el}(V)$ and $\varphi_{mag}(\vec{A})$ are the electrostatic and magnetic contributions to the phase shift, respectively, $\nu$ is the relativistic electron speed and $\text{d}z$ and $\text{d}\vec{z}$ are scalar and vector elements of the electron path, respectively. For a pair of such reversed holograms, half of the sum and half of the difference of the measured phase images represent $\varphi_{el}$ and $\varphi_{mag}$, respectively. Unfortunately, physically turning the sample over by 180°$\degree$ introduces misalignments between the phase images. These misalignments were corrected by using an automated Python script to identify matching features at the sample edges (Walt et al., 2014; Peña et al., 2019; Paterson et al., 2020). Linear distortions caused by the misalignment of the electron optical system of the TEM were corrected by using an affine transformation to provide an optimal match between the edges of each phase image pair, as shown in Fig. 4c. In the present study, the images were found to be elongated by 2% in a direction that forms a 75°$\degree$ angle with the tilt axis by tilting a calibration sample by 180°$\degree$ and tracking fiducial markers. By applying the affine transformation, phase separation artefacts were reduced to below the noise level, such that they were not detectable in line profiles. This approach ensures a consistent separation of $\varphi_{el}$ and $\varphi_{mag}$, as shown in the inset to Fig. 4b. All necessary software packages are listed in Supplementary Material [S3], and examples of intermediate steps of the hologram-to-phase calculation are shown in Supplementary Material [S4].
@@ 2.3 Alignment of phase images
As each $\varphi_{mag}$ image can have non-rectangular borders or contain pixels where the hologram-to-phase calculation resulted in errors, a confidence mask was used to identify the areas in each $\varphi_{mag}$ image that contained reliable measurements.
Each $\varphi_{el}$ image was first flattened to remove background ramps resulting from effects such as electron-beam-induced charging and the perturbed reference wave. A threshold was then used to define a mask that identified where the sample is located at each tilt angle. If elemental mapping was performed, as shown in Fig. 3e, then the threshold could be fine-tuned to include only ferromagnetic material in the mask. These binary masks were used for aligning the phase images in each tilt series. Since the present phase images have a spatial resolution of no better than 6 nm due to the chosen hologram fringe spacing, tomographic reconstruction was found to be highly sensitive to small image misalignments. Furthermore, the sample orientation was not known precisely during acquisition, as only relative tilt angles between phase images $\Delta\theta_{i}$ were measured. The projection positions shown in Fig. 5a were therefore refined by using an axial-symmetry-based implementation of the common lines method (Penczek et al., 1996). As illustrated in Fig. 5b, at every sample tilt angle $\Delta\theta_{i}$ there is a unique projection orientation $\alpha_{i}$, and the two are related via geometrical parameters: $\phi$ is the observed detector rotation relative to the sample tilt axis; $\theta_{0}$ is the starting tilt of the sample; $\alpha_{0}$ is the sample orientation at 0°$\degree$ tilt. $\alpha_{i}$ can be measured by finding the sample symmetry axis from the electrostatic phase maps, as shown in Fig. 5c. The equation
@@
Figure 5: Alignment of tilt series of phase images and generation of a 3D geometrical model of the sample. (a) Schematic diagram showing the angular position and sample outline of each phase image used in the reconstruction. The direction of the electron beam is shown in red. The arcs of rotation of the sample are shown in black. (b) Rotations that define the projections for a conically-symmetrical element, showing the sample tilt $\Delta\theta_{i}$, the projection direction $\alpha_{i}$ and the rotation between the detector and the tilt axis $\phi$. (c) The centre of an axially-symmetrical part of the sample can be identified in each projection and used to align the projections. (d) 3D tomographic reconstruction of the masks in the aligned projections performed to generate a model representing the sample geometry. A 10.2 nm voxel size was used because features smaller than 10 x 10 x 10 nm cannot be resolved in the present study, in part due to tomographic misalignment. The use of a smaller voxel size would not improve the spatial resolution.
@@ 3.1 Uniform region reconstruction
As described in the preceding Section, EH-VFT and MBIR were used to record and reconstruct the 3D distribution of $\vec{M}_{rec}$ in a sample comprising two intersecting nanowires. Fig. 3e shows that the nanowires contain a mixture of cobalt and organic compounds, which transitions from a 60% cobalt content in the core to a pure carbon shell at the surface. Since the cobalt is dispersed within the carbon matrix in nanocrystalline form (Pablo-Navarro et al., 2018) and bulk cobalt has $\mu_{0}M_{s}~{}=~{}1.75$ T (Pablo-Navarro et al., 2018), the nanowires should exhibit $\mu_{0}M_{s}~{}<~{}1.1$ T, as observed for the nanowire distribution in Fig. 10. The few outliers with $\mu_{0}M_{s}~{}>~{}1.1$ T are likely to be artefacts at the interface between the magnetic material and the vacuum region, where magnetic voxels that are designated as non-magnetic force the reconstruction to increase the value of $M_{s}$ locally. In other parts of the sample, surface voxels contain non-magnetic material where $M_{s}$ is close to zero. Surface reconstruction artefacts are present in part because missing wedge correction for the geometrical model is not perfect, as discussed in the Supplementary Material [S1]. The threshold for creating the 3D geometrical model shown in Fig. 5d was refined after the first reconstruction, so that similar amounts of over- and underestimating surface artefacts were present. A discussion of the threshold for the geometrical model and a line trace showing surface artefacts are included in Supplementary Material [S6]. Such surface artefacts are also identifiable for the full sample as voxels with $\mu_{0}M_{s}~{}>~{}1.75$ T, as shown in the histogram in Fig. 10. Such errors could be reduced in the future by using smaller voxels and improved tomographic alignment and distortion removal algorithms. Both nanowires have single magnetic domain configurations at a distance of 200 nm from the intersection, with $\mu_{0}M_{s}~{}=~{}0.7\pm 0.2$ T in the left nanowire and $\mu_{0}M_{s}~{}=~{}0.5\pm 0.2$ T in the right nanowire. $M_{s}$ has previously been measured for varying cobalt contents of FEBID nanowires (Teresa et al., 2016; Pablo-Navarro et al., 2018), but the measurements were corrected to exclude the non-magnetic shell. The latter correction can be replicated by assuming that in Fig. 3e the compositional line profile corresponds to a circular-cross-section nanowire with a 100-nm-diameter magnetic core, resulting in an estimated average cobalt content of 61% with 4% standard deviation, which would correspond to $\mu_{0}M_{s}~{}=~{}0.5\pm 0.1$ T. The reconstructed value of $\vec{M}_{rec}$ is consistent with the predicted values, but shows a higher standard deviation due to surface artefacts.